OpenStudy (anonymous):

Earthquakes release seismic waves that occur in concentric circles from the epicenter of the earthquake. Suppose a seismograph station determines the epicenter of an earthquake is located 9 kilometers from the station. If the epicenter is located at the origin, write the equation for the circular wave that passes through the station.

OpenStudy (anonymous):

if the center of the circle is at the origin and the station is 9 kilometers from the center that means that the radius is 9 and you should easily be able to determine the equation of a circle with the given informaion

OpenStudy (anonymous):

A x2 + y2 = 81 B x2 + y2 = 9 C (x - 9)2 + (y - 9)2 = 0 D (x + 9)2 + (y + 9)2 = 0

OpenStudy (anonymous):

you already know (from the previous question) that the equation of a circle is \[(x - h)^{2} = (y - k)^{2} = r^{2}\] it said that the center is at the origin (If the epicenter is located at the origin...) so (h, k) = (0, 0). since the station is 9 km away, the radius (r) will be 9. plug in those values and you should have it.

OpenStudy (anonymous):

I do online school so its really hard :( to learn math on my own and i have 17 days left to finish everything

OpenStudy (anonymous):

the previous question has nothing to do with this question

OpenStudy (anonymous):

circle equation \[(x-x_0)^2+(y-y_0)^2=r^2\] where (xo,yo) is the location of the center of the circle and r is the radius

OpenStudy (anonymous):

So A?

OpenStudy (anonymous):

i already gave you everything you need to solve the problem so have more confidence in your own abilities

OpenStudy (anonymous):

but yes its right

OpenStudy (anonymous):

Right the radius is 9 so 9 squared is 81?

OpenStudy (anonymous):

okay thank you soooo much :DD